OpenStudy (anonymous):

The length of an airplane hanger is 20ft less than 4 times its height h. The width of the hanger is 10 ft more than 2 times its height. What polynomial expression represents the volume of the hanger?

OpenStudy (anonymous):

Length=4h-20 Width=2h+10 Volume of hanger=L*W*H h(4h-20)(2h+10)

OpenStudy (anonymous):

(4h^2-20h)(2h+10)

OpenStudy (anonymous):

8h^3+40h^2 -40h^2-30h

OpenStudy (anonymous):

8h^3-30h=final answer I think

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

A commented solution using Mathematica is attached.

OpenStudy (anonymous):

Thanks